# Multi-modal Evidential Fusion Network for Trusted PET/CT Tumor Segmentation

### v1

@@ 4.4.1 Effectiveness of each component
| Dataset | Methods | DBTF | MTGA | DFC | DSC(%) | Jaccard(%) | HD95(mm) | Sens(%) | Pre(%) |
@@
| Baseline+DBTF | - | - | 81.63 25.17 | 74.43 26.31 | 4.68 10.81 | 83.31 22.41 | 88.18 20.59 | ||
| Baseline+DBTF+MTGA | - | 82.29 25.46 | 75.22 26.45 | 3.05 7.91 | 86.17 19.95 | 86.29 22.79 | |||
| Baseline+DBTF+DFC | - | 82.43 24.29 | 75.28 25.64 | 3.05 7.91 | 85.07 21.16 | 87.33 20.71 | |||
@@
| Baseline+DBTF | - | - | 81.63 25.71 | 71.22 29.30 | 3.52 7.09 | 77.19 30.92 | 89.43 15.84 | ||
| Baseline+DBTF+MTGA | - | 82.45 23.02 | 69.49 28.60 | 3.62 7.32 | 74.38 30.03 | 89.67 15.63 | |||
| Baseline+DBTF+DFC | - | 83.13 23.55 | 71.44 26.31 | 3.34 8.29 | 73.31 22.41 | 88.18 20.59 | |||
@@
To better perform multi-modal segmentation, DBTF, MTGA and DFC modules were incorporated in our segmentation model. To validate the effectiveness of each component, ablation experiments are conducted. The baseline network for comparison is the UNet without the DBTF, MTGA and DFC modules. In Baseline, PET features and CT features are fused simply by concatenation and input to to obtain the final result. The Baseline+DBTF is the one by incorporating the DBTF module to the Baseline model. The Baseline+DBTF+MTGA is the model by further incorporating MTGA to Baseline+DBTF. The Baseline+DBTF+DFC denotes the model by incorporating DFC to Baseline+DBTF. Finally, upon incorporating all three components into the Baseline we obtain our proposed MEFN. Ablative experiments were conducted respectively on the AutoPET and Hecktor datasets whose results are shown in Tabel 2. From this table, it can be observed that on the AutoPET dataset, the result of Baseline+DBTF is significantly superior to that of the Baseline model (81.63 vs. 79.91 in DSC), which indicates that the DBTF module effectively utilizes the uncertainties of PET features, CT features and fused features to achieve more reasonable fusion of segmentation results. After incorporating MTGA, the model achieves a noticeable improvement in HD95 and sensitivity, which indicates that MTGA can effectively assist the network in Tumor-Related initialization and hence aids in identifying more whole-body tumors. With the introduction of DFC, the result is also improved compared to Baseline+DBTF (82.43 vs. 81.63 in DSC), demonstrating that the proposed feature fusion module effectively eliminates redundant features and improves segmentation performance. After incorporating all components, further improvements can be observed, which indicates that all modules contribute corporatively to the proposed method.
@@ 4.4.2 Effectiveness of Uncertainty Calibrator
In this work, the Uncertainty Calibrator (UC) is proposed to achieve semantic correction of decoder features during feature fusion. To validate the effectiveness of the proposed method, we introduced perturbations of Gaussian noise at different levels (with noise variance ) and patch-size random masking (mask ratio ) into the AutoPET and Hecktor datasets, respectively. For patch-size random masking, we use 33 black boxes to cover the original input image with a uniform distribution, and the mask ratio represents the proportion of black boxes to the total number of pixels. We compare the performance of our MEFN, MEFN without UC (referred to as w/o. UC) and other models on out-of-distribution data (with large uncertainty), whose results are shown in Fig. 6. From this figure, it can be observed that as the perturbation level increases, the performance of other models deteriorate drastically, while our proposed MEFN performs much more robust to random noise and random masking. We notice that after removing UC, the performance degradation rate of our model without UC is comparable to CML (Zhang et al., 2021) (blue line and green line), which indicates that the model loses the ability to calibrate semantic features in large uncertainty regions after removing UC from our MEFN model.
@@ 4.4.3 Effectiveness of Uncertainty Perceptual Loss
In our work, the Uncertainty Perceptual Loss () is proposed to provide supervision for uncertainty. To validate the effectiveness of in model training, we compare performance of three models trained respectively under the combination of the baseline loss and other three different loss functions used for supervising uncertainty: uncertainty loss (Huang et al., 2021b) (), calibrated uncertainty loss (Zou et al., 2023) () and a variant of calibrated uncertainty loss (Zou et al., 2023) (). Table 3 gives the segmentation performance of models trained under different loss functions. On the AutoPET dataset, compared to the model trained under the baseline loss , models trained under (Huang et al., 2021b), (Zou et al., 2023) or (Zou et al., 2023) yield slight improvements in segmentation performance (77.92 vs. 78.65 vs. 77.93 vs. 78.46 in DSC). Specifically, when incorporating our proposed Uncertainty Perceptual Loss, the overall segmentation performance is significantly improved compared to that of the one trained under the baseline loss (77.92 vs. 82.45 in DSC). On the Hecktor dataset, although the model trained under does not perform the best in terms of HD95 and precision, there is still a substantial improvement in overall metrics compared to the one trained under the baseline loss . Fig. 7 illustrates changes in DSC with iterations during the training processes of models under different loss functions. It can be observed from Fig. 7 that the DSC of the model trained under shows a decreasing trend with increased iterations, indicating overfitting. After incorporating to the loss function, not only does the overall segmentation performance improve, but the overfitting problem in the later training stages is also alleviated. With the incorporation of (Huang et al., 2021b), the model exhibits strong oscillations in the early training stages, which is possibly due to inaccurate quantification of uncertainty by the model in initial stages. However, upon introducing an annealing factor , our proposed stabilizes the early training of the model.
@@ 4.4.4 Effectiveness of CFL
In our work, a novel task-driven initialization module called CFL is introduced for multi-modal segmentation tasks. Here, the effectiveness of such an initialization method for multi-modal segmentation is evaluated in terms of the test error defined as follows:
where , and are as defined in Section 4.2, (true negatives) represents the number of negative samples correctly classified as negative. We compare four different initialization methods in terms of test error curves during model training. These four initialization methods considered are Normal (standard normal distribution), Xavier (Glorot and Bengio, 2010), Kaiming (He et al., 2015) and Orthogonal (Ahmad et al., 2023). Fig. 8 illustrates curves of the test error on the AutoPET dataset, while Table 4 provides the final test error for both AutoPET and Hecktor datasets.
As a baseline, the Normal initialization strategy achieves final test errors of 22.13 and 17.72 on the AutoPET and Hecktor datasets, respectively. When using CFL initialization method, test error on the same dataset are reduced to 17.62 and 15.76, respectively, which demonstrates a significant improvement. Conversely, other initialization methods show minor improvements or even performance degradation. From the Fig. 8, it can be clearly seen that when using Orthogonal (Saxe et al., 2013) initialization method, the model have fallen into a local optimum, resulting in poor generalization performance. It is evident that curves for the test error when utilizing the CFL initialization method are notably lower than those with other initialization methods and exhibit smoother trends. These results convincingly demonstrate the effectiveness of CFL.
@@ 4.5 Visual assessment on segmentation uncertainty
Intuitionally, the uncertainty output from different models should be positively correlated to data perturbation (i.e., the output uncertainty should increase with the incresement of perturbation levels of input data). The conservation of such a positive correlation is a measure of the confidential level of the output uncertainty of different models. Thus, we introduce perturbation of different levels into the dataset with the same perturbation methods as described in Section 6 and then conduct segmentation with different models. The correlation of the output uncertainty with the level of perturbation is employed to assess the reliability of the output uncertainty. We visually compare the uncertainty output from the DBTF module of our MEFN with that output from MMEF (Huang et al., 2021a) for the segmentation results. Fig. 9 gives the segmentation results of two sample images from the AutoPET dataset along with their segmentation uncertainty. The top part of this figure correspond to the first sample while the bottom part correspond to the second sample, with columns from left to right corresponding to sample CT image, sample PET image, ground truth label, segmentation results from the MMEF and their uncertainty, segmentation results from our MEFN and their uncertainty. Rows of this figure correspond to the sample images themselves and and their perturbed versions. Tumor regions are marked with red boxes. It can be evidently seen from the 4th and 6th columns of both samples that the segmentation results from both models become inferiors to different extent as the level of perturbation and the degree of Out-Of-Distribution (OOD) data increase. Concurrently, the uncertainty predicted by our MEFN also consistently increases (as shown in the 7th columns in both samples), which implies a positive correlation between the uncertainty and the level of perturbation. In contrast, the uncertainty predicted with MMEF (Huang et al., 2021a) inconsistently decreases (as shown in the 5th columns in both samples) in some cases, which obviously violates the positive correlation requirement and is probably due to overconfidence in the model. These observations indicate that our proposed MEFN avoids blind confidence and provides radiologists and doctors with more credible uncertainty of the segmentation results. Moreover, beside tumor regions, MMEF (Huang et al., 2021a) exhibits significant uncertainty in non-tumor regions (as pointed by red arrows in the uncertainty images in the third row in the first sample and first, seconde and third rows in the second sample). In contrast, MEFN only shows uncertainty in tumor regions(as shown in the 7th columns in both samples), indicating precise localization ability for tumors and more accurate handling of uncertainty in non-tumor regions. Furthermore, as shown in the 5th and 7th columns of the first rows in both samples, our MEFN shows less uncertainty as compared with MMEF (Huang et al., 2021a), which indicates that our MEFN can provide more reliable segmentation results.
@@ 5 Conclusion
This paper proposes a novel multi-modal Evidential fusion network for PET/CT tumor segmentation. The motivation of this work is to develop a method that can fuse the somewhat inconsistent functional and anatomical information for joint segmentation of tumors in both the CT and PET images. In this method, a Generative Adversarial Network-based Cross-modali-ty Feature Learning module is first designed to obtain task-driven initialization of network parameters by reducing the domain gap. Then, two U-Net networks are utilized to extract CT and PET features for feature fusion, during which a Dual-attention Calibrating module is specifically designed to reduce the impact of redundancy and confliction between features of different modalities on segmentation tasks. Finally, the output results are trustworthily fused based on the Dempster-Shafer Theory according to uncertainty. Extensive experiments were performed on two publicly available datasets whose results dem-onstrate the superiority of the proposed MEFN method over other state-of-the-art multi-modal image segmentation methods. The improvements achieved with our MEFN method over other state-of-the-art methods in PET/CT multi-modal tumor segmentation amount to 2.15 and 3.23 in DSC on the AutoPET dataset and the Hecktor dataset, respectively.
@@
Yuxuan Qi: Conceptualization, Writing - review and editing, Writing - original draft, Visualization, Software, Methodology, Investigation, Data curation, Formal Analysis. Li Lin: Conceptualization, Writing - review and editing, Methodology, Investigation, Supervision, Validation, Formal Analysis. Jingya Zhang: Resource, Data curation, Funding acquisition. Jiajun Wang: Conceptualization, Writing - review and editing, Validation, Supervision, Project administration, Funding acquisition, Conceptualization. Bin Zhang: Resource, Investigation, Data curation, Funding acquisition.

### v2

@@ 4.4.1 Effectiveness of each component
Dataset | Methods | DTF | TGA | DFC | DSC(%) | Jaccard(%) | HD95(mm) | Sens(%) | Pre(%) |
@@
Baseline+DTF | - | - | 81.63 25.17 | 74.43 26.31 | 4.68 10.81 | 83.31 22.41 | 88.18 20.59 | ||
Baseline+DTF+TGA | - | 82.29 25.46 | 75.22 26.45 | 3.05 7.91 | 86.17 19.95 | 86.29 22.79 | |||
Baseline+DTF+DFC | - | 82.43 24.29 | 75.28 25.64 | 3.05 7.91 | 85.07 21.16 | 87.33 20.71 | |||
@@
Baseline+DTF | - | - | 81.63 25.71 | 71.22 29.30 | 3.52 7.09 | 77.19 30.92 | 89.43 15.84 | ||
Baseline+DTF+TGA | - | 82.45 23.02 | 69.49 28.60 | 3.62 7.32 | 74.38 30.03 | 89.67 15.63 | |||
Baseline+DTF+DFC | - | 83.13 23.55 | 71.44 26.31 | 3.34 8.29 | 73.31 22.41 | 88.18 20.59 | |||
@@
To better perform multi-modal segmentation, DTF, TGA and DFC modules are incorporated in our segmentation model. To evaluate the effectiveness of each component, ablation experiments are conducted on the AutoPET and Hecktor datasets, with the results presented in Table 2. The baseline network for comparison is the UNet without the DTF, TGA and DFC modules. In Baseline, PET features and CT features are fused simply by concatenation and input to to obtain the final result. It is worth noting that baseline still contains CFL to align features across different modalities and the outputs of CT branch and PET branch are used for deep supervision to constrain the fusion result. Baseline+DTF incorporates the DTF module into the Baseline model. Baseline+DTF+TGA further adds TGA, while Baseline+DTF+DFC integrates DFC into Baseline+DTF. Finally, upon incorporating all three components into the Baseline we obtain our proposed MEFN. Notably, as shown in Table 2, our baseline surpasses several state-of-the-art methods. This may be attributed to our use of CFL to learn more robust feature representations and the application of results from the CT and PET branches as deep supervision. On the AutoPET dataset, the result of Baseline+DTF is significantly superior to that of the Baseline model (81.63 vs. 79.91 in DSC), which indicates that the DTF module effectively utilizes the uncertainties of PET result, CT result and fusion result to achieve more reasonable weight allocation in the decision-level fusion. After incorporating TGA, the network focuses more on the tumor regions, making the model achieve a noticeable improvement in HD95 and sensitivity. With the introduction of DFC, the result is improved compared to Baseline+DTF (82.43 vs. 81.63 in DSC). Such improvement may stem from the mutual attention mechanism of CMA and the uncertainty calibration mechanism of UC which effectively reduce the redundant features and conflict features in the training process of the network. After incorporating all components, further improvements can be observed, which indicates that all modules contribute corporatively to the proposed method.
@@ 4.4.2 Effectiveness of Uncertainty Calibrator
In this work, the Uncertainty Calibrator (UC) is proposed to achieve semantic correction of decoder features during feature fusion. To validate the effectiveness of the proposed method, we introduced perturbations of Gaussian noise at different levels (with noise variance ) and patch-size random masking (mask ratio ) into the AutoPET and Hecktor datasets, respectively. For patch-size random masking, we use 33 black boxes to cover the original input image with a uniform distribution, and the mask ratio represents the proportion of black boxes to the total number of pixels. We compare the performance of our MEFN, MEFN without UC (referred to as w/o. UC) and other models on out-of-distribution data (with large uncertainty), whose results are shown in Fig. 6. From this figure, it can be observed that as the perturbation level increases, the performance of other models deteriorate drastically, while our proposed MEFN performs much more robust to random noise and random masking. We notice that after removing UC, the performance degradation rate of our model without UC is comparable to CML [38] (blue line and green line), which indicates that the model loses the ability to calibrate semantic features in large uncertainty regions after removing UC from our MEFN model.
@@ 4.4.3 Effectiveness of Uncertainty Perceptual Loss
In our work, the Uncertainty Perceptual Loss () is proposed to provide supervision for uncertainty. To validate the effectiveness of in model training, we compare performance of three models trained respectively under the combination of the baseline loss and other three different loss functions used for supervising uncertainty: uncertainty loss [42] (), calibrated uncertainty loss [43] () and a variant of calibrated uncertainty loss [43] (). Table 3 gives the segmentation performance of models trained under different loss functions. On the AutoPET dataset, compared to the model trained under the baseline loss , models trained under [42], [43] or [43] yield slight improvements in segmentation performance (77.92 vs. 78.65 vs. 77.93 vs. 78.46 in DSC). Specifically, when incorporating our proposed Uncertainty Perceptual Loss, the overall segmentation performance is significantly improved compared to that of the one trained under the baseline loss (77.92 vs. 82.45 in DSC). On the Hecktor dataset, although the model trained under does not perform the best in terms of HD95 and precision, there is still a substantial improvement in overall metrics compared to the one trained under the baseline loss . Fig. 7 illustrates changes in DSC with iterations during the training processes of models under different loss functions. It can be observed from Fig. 7 that the DSC of the model trained under shows a decreasing trend with increased iterations, indicating overfitting. After incorporating to the loss function, not only does the overall segmentation performance improve, but the overfitting problem in the later training stages is also alleviated. With the incorporation of [42], the model exhibits strong oscillations in the early training stages, which is possibly due to inaccurate quantification of uncertainty by the model in initial stages. However, upon introducing an annealing factor , our proposed stabilizes the early training of the model.
@@ 4.4.4 Effectiveness of CFL
In our work, a novel modality translation stage, called CFL, is introduced for alternate training and provides initialization parameters for the model in the MTF stage. To validate the effectiveness of CFL, we compare our method with four common initialization methods that do not use alternating training, and evaluate the model’s training progress and performance using DSC. These four initialization methods considered are Random (standard normal distribution), Xavier [44], Kaiming [45] and Orthogonal [30]. Fig. 8 illustrates curves of the DSC on the AutoPET dataset, while Table 4 provides the final DSC for both AutoPET and Hecktor datasets.
As a baseline, the Random initialization strategy achieves final DSC of 77.87 and 82.28 on the AutoPET and Hecktor datasets, respectively. When using CFL, DSC on the same dataset are improved to 82.45 and 83.35, respectively, which demonstrates a significant improvement. Conversely, other initialization methods show minor improvements or even performance degradation. From the Fig. 8, it can be clearly seen that when using Orthogonal [46] initialization method, the model have fallen into a local optimum, resulting in poor generalization performance. It is evident that curves for the DSC when utilizing the CFL stage are notably higher than those with other initialization methods and exhibit smoother trends. These results convincingly demonstrate the effectiveness of CFL stage.
@@ 4.5 Visual assessment on segmentation uncertainty
Intuitionally, the uncertainty output from different models should be positively correlated to data perturbation (i.e., the output uncertainty should increase with the incresement of perturbation levels of input data). The conservation of such a positive correlation is a measure of the confidential level of the output uncertainty of different models. Thus, we introduce perturbation of different levels into the dataset with the same perturbation methods as described in Section 6 and then conduct segmentation with different models. The correlation of the output uncertainty with the level of perturbation is employed to assess the reliability of the output uncertainty. We visually compare the uncertainty output from the DTF module of our MEFN with that output from MMEF [31] for the segmentation results. Fig. 9 gives the segmentation results of two sample images from the AutoPET dataset along with their segmentation uncertainty. The top part of this figure correspond to the first sample while the bottom part correspond to the second sample, with columns from left to right corresponding to sample CT image, sample PET image, ground truth label, segmentation results from the MMEF and their uncertainty, segmentation results from our MEFN and their uncertainty. Rows of this figure correspond to the sample images themselves and and their perturbed versions. Tumor regions are marked with red boxes. It can be evidently seen from the 4th and 6th columns of both samples that the segmentation results from both models become inferiors to different extent as the level of perturbation and the degree of Out-Of-Distribution (OOD) data increase. Concurrently, the uncertainty predicted by our MEFN also consistently increases (as shown in the 7th columns in both samples), which implies a positive correlation between the uncertainty and the level of perturbation. In contrast, the uncertainty predicted with MMEF [31] inconsistently decreases (as shown in the 5th columns in both samples) in some cases, which obviously violates the positive correlation requirement and is probably due to overconfidence in the model. These observations indicate that our proposed MEFN avoids blind confidence and provides radiologists and doctors with more credible uncertainty of the segmentation results. Moreover, beside tumor regions, MMEF [31] exhibits significant uncertainty in non-tumor regions (as pointed by red arrows in the uncertainty images in the third row in the first sample and first, seconde and third rows in the second sample). In contrast, MEFN only shows uncertainty in tumor regions(as shown in the 7th columns in both samples), indicating precise localization ability for tumors and more accurate handling of uncertainty in non-tumor regions. Furthermore, as shown in the 5th and 7th columns of the first rows in both samples, our MEFN shows less uncertainty as compared with MMEF [31], which indicates that our MEFN can provide more reliable segmentation results.
@@ 5 Conclusion
This paper proposes a novel multi-modal Evidential fusion network for PET/CT tumor segmentation. The motivation of this work is to develop a method that can efficient and accurate fuse PET metabolic information and CT anatomical information. In this method, a GAN-based Cross-modality Feature Learning stage is first designed to align features across different modalities by mapping them to a shared feature space, thereby learning more robust feature representations. Then, two U-shaped networks are utilized to extract CT and PET features for feature fusion, during which a Dual-attention Calibrating module is specifically designed to reduce the impact of redundancy and confliction between features of different modalities on segmentation tasks. Finally, the output results are trustworthily fused based on the Dempster-Shafer Theory according to uncertainty. Extensive experiments are performed on two publicly available datasets whose results demonstrate the superiority of the proposed MEFN method outperforms other state-of-the-art multi-modal image segmentation methods. The improvements achieved with our MEFN method over other state-of-the-art methods in PET/CT multi-modal tumor segmentation amount to 3.10 and 3.23 in DSC on the AutoPET dataset and the Hecktor dataset, respectively.
@@
Yuxuan Qi: Conceptualization, Writing - review and editing, Writing - original draft, Visualization, Software, Methodology, Investigation, Data curation, Formal Analysis. Li Lin: Conceptualization, Writing - review and editing, Methodology, Investigation, Supervision, Validation, Formal Analysis. Jingya Zhang: Resource, Data curation, Funding acquisition. Jiajun Wang: Conceptualization, Writing - review and editing, Validation, Supervision, Project administration, Funding acquisition. Bin Zhang: Resource, Investigation, Data curation, Funding acquisition.